# leftmargin=10pt, rightmargin=10pt, vskip=0pt

### v1

@@ A.2.6 UTR-LM
The preprocessing of 5’ UTR sequences for UTR-LM involved a 4-step pipeline: First, all coding sequence (CDS) and non-5’ UTR fragments were removed from the raw sequences. Second, duplicate sequences were identified and removed. Third, the sequences were truncated to fit within a range of 30 to 1022 base pairs. Finally, incorrect and low-quality sequences were filtered out.
A.2.7 BEACON-B
@@
The pre-training configs of BEACON-B and BEACON-B512 are shown as Tab 9 and Tab 10
Table 9: Configuration settings for the BEACON-B pre-training
@@
Table 10: Configuration settings for the BEACON-B512 pre-training
@@ A.3 Computational Resources
Table 11: The number of Top2 performance among different tokenizers and positional encodings
@@ A.4 Additional results
For the experiments 4 on component analysis of the baseline RNA language model, we further counted the number of top performances for different tokenizers and positional encoding as shown in Tab 11. The effectiveness of Single nucleotide tokenizer and ALiBi can be demonstrated directly.
A.5 Broader Societary Impacts
@@
This work is dedicated to establishing a robust and versatile benchmark for RNA-related tasks, enhancing the understanding of RNA sequences across diverse applications. Our benchmark, encompassing a variety of RNA tasks, aims to rigorously evaluate the efficacy of different RNA representation covering structural analysis, functional studies, and engineering applications. By doing so, it provides a critical assessment of their potential utility in real-world scenarios, thereby laying a foundational framework for applying deep learning in fields such as medical research and genetics.
@@
However, it is also important to acknowledge the dual-use nature of any powerful technology, including those developed from our benchmark. For instance, the enhanced ability to manipulate RNA sequences might be misused, such as in the creation of adverse viral agents. Moving forward, it is crucial to address these risks. We will develop and implement guidelines for the ethical and safe use of our benchmark in the future, ensuring that it contributes positively to society and does not enable harmful applications.
A.6 Assets
A.6.1 Software and Libraries
The open-source software, and corresponding licenses are presented in Tab. 12. The data, licenses and corresponding URL are presented in Tab. 13.

### v2

@@ A.3.6 UTR-LM
The preprocessing of 5’ UTR sequences for UTR-LM involved a 4-step pipeline: First, all coding sequences (CDS) and non-5’ UTR fragments were removed from the raw sequences. Second, duplicate sequences were identified and removed. Third, the sequences were truncated to fit within a range of 30 to 1022 base pairs. Finally, incorrect and low-quality sequences were filtered out.
A.3.7 BEACON-B
@@
The pre-training configs of BEACON-B and BEACON-B512 are shown as Table 14 and Table 15.
Table 14: Configuration settings for the BEACON-B pre-training.
@@
Table 15: Configuration settings for the BEACON-B512 pre-training.
@@ A.4 Computational Resources
A.5 Computational Complexity Comparisons
We have compared the currently available pre-training resource consumption of RNA language models in Table 5, shown via GPU Days (Days * GPUs). We also compared the FLOPs and MACs of the different models in Table 16. The detailed computational costs of the models are outlined as follows (following the sequence lengths at which the models were pretrained). Despite our model employing a vanilla BERT-base model with 12 layers, which doesn’t minimize FLOPs, it still manages to achieve the lowest resource usage, requiring only 3.58 GPU Days. This efficiency is primarily due to our use of a modestly sized, filtered ncRNA dataset combined with a lightweight pre-training schedule (80K steps and a batch size of 512).
Table 16: FLOPs, MACs, and GPU days for various models.
Method FLOPs (G) MACs (G) GPU Days (Days * GPUs)
RNA-FM 233.88 116.77 30 * 8 = 240
RNABERT 0.91 0.46 -
RNA-MSM 201.46 84.56 -
SpliceBERT-H510 19.27 9.63 7 * 8 = 56
SpliceBERT-MS510 19.27 9.63 7 * 8 = 56
SpliceBERT-MS1024 38.69 19.33 7 * 8 + 3 * 4 = 68
UTR-LM-MRL 5.76 2.83 -
UTR-LM-TE&EL 5.76 2.83 -
UTRBERT-3mer 511.41 255.56 38 * 4 = 152
UTRBERT-4mer 511.41 255.56 38 * 4 = 152
UTRBERT-5mer 511.41 255.56 38 * 4 = 152
UTRBERT-6mer 511.41 255.56 38 * 4 = 152
BEACON-B 198.48 99.13 1.3 * 8 = 10.4
BEACON-B512 87.02 43.49 0.895 * 4 = 3.58
Table 17: The number of Top2 performance among different tokenizers and positional encodings.
@@ A.6 Additional Results
For the experiments 4 on component analysis of the baseline RNA language model, we further counted the number of top performances for different tokenizers and positional encoding as shown in Table 17. The effectiveness of Single nucleotide tokenizer and ALiBi can be demonstrated directly.
In addition, we further collected additional testing datasets to evaluate BEACON-B and other models for their generalization. On the one hand, we collected the contact map data W19 from Rfam [112], results with direct generalization are outlined in Table 18. It shows our BEACON-B models can achieve significant performance.
Table 18: Additional Results on W19 Dataset of CMP.
Method Metric (SC %)
PLMC 48.00
RNAcontact 72.00
RNA-FM 55.98
RNABERT 26.19
SpliceBERT-H510 66.00
UTR-LM-MRL 60.00
UTRBERT-3mer 22.27
UTRBERT-6mer 53.41
BEACON-B 76.01
BEACON-B512 73.56
On the other hand, we collected two unseen APA isoform libraries HSPE1 and WHAMMP2 from MPRAs [94, 9], results with direct generalization are outlined in Table 19. On the task of predicting 3’UTR functions, BEACON-B’s generalization capabilities show potential to rival those of UTRBERT-3mer, which was specifically pre-trained on 3’UTR sequences.
Table 19: Additional Results on HSPE1 and WHAMMP2 Dataset of APA.
Method HSPE1 Metric ($R^{2}$) WHAMMP2 Metric ($R^{2}$)
APARENT 32.95 18.56
RNA-FM 35.35 16.43
RNABERT 19.04 3.43
SpliceBERT-H510 17.58 4.81
UTR-LM-MRL 28.62 8.73
UTRBERT-3mer 33.73 17.56
UTRBERT-6mer 42.71 20.17
BEACON-B 29.98 15.35
BEACON-B512 33.22 17.83
A.7 Ablation Study
We carried out detailed ablation experiments in the pre-training setting in Table 20. We used the same pre-training setup of BEACON-B512 (pre-training on 512-length sequences of human-filtered ncRNA and using FlashAttn). This setup allowed us to scrutinize the impact of various tokenizers and positional encodings. For downstream evaluation, in order to minimize computational resources and accelerate experimental procedures, we evaluated three major RNA task categories: Distance Map Prediction (DMP) from structural tasks, APA Isoform Prediction (APA) from functional tasks, and Vaccine Degradation Prediction (VDP) from engineering tasks.
Table 20: Performance comparison of different methods on CMP, APA, and VDP tasks.
Method Tokenizer Positional Encoding Metric (P@L %) Metric ($R^{2}$) Metric (MCRMSE↓)
BPE-ALiBi BPE ALiBi 49.08 64.98 0.640
Non-overlap-ALiBi Non-overlap ALiBi 47.68 66.61 0.639
6mer-ALiBi 6mer ALiBi 48.79 70.77 0.372
Single-RoPE Single RoPE 12.84 33.86 0.462
Single-APE Single APE 50.76 66.20 0.327
BEACON-B512 Single ALiBi 56.82 72.00 0.320
A.8 More Discussion of Future Work and Limitations
We provide additional discussions with the following two aspects:
RNA tasks with other inputs: To ensure fairness in comparison, we standardize input formats by using RNA sequences across all tasks and all models. This approach was necessary to maintain consistency in evaluations. However, it may not fully capture the potential of models that are designed to leverage richer input data, such as the scRNA-seq data for cell annotation and gene regulation network which uses genes with expression level [105] and inverse RNA folding using structure as input [46]. Future work could explore alternative evaluation settings that allow for the use of more complex inputs, thereby providing a more nuanced assessment of model performance.
In particular, although the task of inverse RNA folding extends beyond the current scope of this benchmark, we plan to extend our benchmark to support more complex tasks involving diverse types of input, broadening our coverage in RNA research significantly in the future. Specifically, for inverse RNA folding, we will consider datasets and tasks such as Eterna100v1 [4] and Eterna100v2 [59] for standard inverse RNA folding evaluations and PseudoBase++ [106] to evaluate the capability of inverse folding extending to pseudoknots. Regarding models, we will evaluate significant inverse RNA folding methods such as RNAinverse [46], MCTS-RNA [118], LEARNA [86], MetaLEARNA [86], gRNAde [56], and RiboDiffusion [49]. We are also considering influential inverse protein folding models like StructGNN [52], GVP-GNN [55], and ProteinMPNN [27], PiFold [37], given their potential to offer new insights into RNA structure modeling, owing to the inverse task between modeling protein and RNA structures. We could also explore using cross-modal generation methods [2, 64, 81, 80] to bridge the structure and sequence for the inverse task. For evaluation metrics, we could consider assessing sequence similarity, sequence diversity, and structural similarity, using indicators such as recovery rate, diversity, and structural similarity.
Other important RNA sequence tasks: While our current benchmark effectively evaluates RNA language models across a spectrum of tasks, we acknowledge the out-of-selection of certain complex tasks such as 3D RNA structure prediction. This is primarily due to several challenges: (1) The majority of open-source RNA language models have not been developed to perform 3D structure prediction tasks. (2) The acquisition of high-quality structural data, often requiring Multiple Sequence Alignment (MSA) data, is prohibitively expensive. This, coupled with the limited availability of such data, poses a significant barrier [108]. (3) The computational pipeline for predicting 3D structures (e.g., structural modules) is exceedingly complex and resource-intensive, often necessitating days of computation on high-performance GPUs [65]. In future extensions of our benchmark, we aim to incorporate more diverse tasks, including those requiring complex data inputs and significant computational resources, to better encompass the breadth of challenges in RNA research.
A.9 Broader Societary Impacts
@@
This work is dedicated to establishing a robust and versatile benchmark for RNA-related tasks, enhancing the understanding of RNA sequences across diverse applications. Our benchmark, encompassing a variety of RNA tasks, aims to rigorously evaluate the efficacy of different RNA representations covering structural analysis, functional studies, and engineering applications. By doing so, it provides a critical assessment of their potential utility in real-world scenarios, thereby laying a foundational framework for applying deep learning in fields such as medical research and genetics.
@@
However, it is also important to acknowledge the dual-use nature of any powerful technology [48, 122, 31], including those developed from our benchmark. For instance, the enhanced ability to manipulate RNA sequences might be misused, such as in the creation of adverse viral agents. Moving forward, it is crucial to address these risks. We will develop and implement guidelines for the ethical and safe use of our benchmark in the future, ensuring that it contributes positively to society and does not enable harmful applications.
A.10 Assets
@@
A.10.1 Software and Libraries
The open-source software, and corresponding licenses are presented in Table 21. The data, licenses and corresponding URL are presented in Table 22.
@@ A.10.1 Software and Libraries
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See the abstract and last three paragraphs of Sec. 1
    2. (b)
      Did you describe the limitations of your work? [Yes] We discuss the limitations in Sec. 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix A.9
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We provide a URL for all the source data, code and instructions
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Sec 3, Sec 5 and Appendix A.1
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Table 3 and Table 4
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix A.4
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Sec 3 and Sec 4
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix A.10
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We contribute a new database which can be downloaded via the URL
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]